The largest number of molecules is in
a. 36 g of water
b. 28 g of CO
c. 46 g of ethy1 alcohol
d. 54 g of nitrogen pentaoxide (N2O5)

Grade:10

3 Answers

Deepak Patra
9 years ago
Sol. Number of molecules present in 36 g of water = 36/18×N A = 2N A Number of molecules present in 28 g of CO = 28/28×N A = N A Number of molecules present in 46 g of C2H5OH = 46/46×N A = 2N A Number of molecules present in 54 g of N2O5 = 54/108×N A = 0.5N A Here N A is Avogadro’s number. Hence, 36 g of water contain the larges (2N A) number of molecule.
Anjali kumari
5 years ago
As in (a) the number of moles=2 moles.And in (b) the number of moles=1mole.And in (c) the number of moles=2 mole.And in (d) the number of moles=1/2 moles.So largest number of molecules =2 moles.So the answer is (a). 
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Number of molecules present in 36 g of water = 36/18×N A
= 2N A
Number of molecules present in 28 g of CO = 28/28×N A = N A
Number of molecules present in 46 g of C2H5OH = 46/46×N A = 2N
A Number of molecules present in 54 g of N2O5 = 54/108×N A = 0.5N A
Here N A is Avogadro’s number.
Hence, 36 g of water contain the larges (2N A) number of molecule.

Thanks and Regards